# Discriminative Fine-tuning of LVLMs

### v2

@@ Abstract
Contrastively-trained Vision-Language Models (VLMs) like CLIP have become the de facto approach for discriminative vision-language representation learning. However, these models have limited language understanding, often exhibiting a “bag of words” behavior. At the same time, Large Vision-Language Models (LVLMs), which combine vision encoders with LLMs, have been shown capable of detailed vision-language reasoning, yet their autoregressive nature renders them less suitable for discriminative tasks.
@@
footnotetext: * Equal contribution. Listing order is random.
@@ 1 Introduction
Concomitantly, inspired by the success of LLMs [5, 42] in acting as generalist assistants [15], a series of works combine pretrained vision encoders and LLMs [26, 27, 50] to construct Large Vision-Language Models (LVLMs) capable of performing interactive multi-modal conversations between the model and the user. Among others, these models have been shown capable of exhibiting strong reasoning and vision-language understanding capabilities, offering fine-grained and detailed responses [27, 26, 9, 11]. However, they are trained with a next-token prediction loss in an autoregressive manner, which appears less suitable for direct utilization in discriminative image-text tasks (e.g. image-text retrieval).
@@
In our approach, following the (independent) two-towers paradigm, the vision embeddings are produced by passing the image through the entire LVLM, and the text embeddings are produced by passing the text through the LLM of the LVLM. Intuitively, for vision embedding, the LLM acts as an information extractor/processor that refines the visual information extracted by the data while simultaneously aligning the visual representations with the textual ones. We coin our approach VladVA: Vision-Language Adaptation for Discriminative Visual Assistant.
Our main contributions are:
  • We devise a carefully designed optimization framework that utilizes image-text pairs of variable length and granularity for model training (i.e. both short and long captions). Using this data, the model is trained with both contrastive and next-token prediction losses, which are both shown to be necessary for unlocking strong discrimination and compositionality capabilities. Our design choices are accompanied by ablation studies which justify the necessity of our framework’s components.
@@ 2.1 Large Vision Language Models (LVLMs)
Despite their strong generative and comprehension abilities [34], current LVLMs are primarily restricted to generative tasks. Only very recently, Jiang et al. [23], inspired by the recent progress in NLP [4, 24] adapted a LLaVA-NeXT [26] model to discriminative tasks with the help of a contrastive-like loss, using text data only. We note that unlike [23], we introduce a training framework that learns from multi-turn image-text pairs (as opposed to text only) using a novel formulation that jointly combines a contrastive loss with a next-token prediction, reflecting the data characteristics and inducing a gradual representation buildup. Moreover, we compare our approach with E5-V [23], significantly improving upon their results despite using smaller/lighter models.
@@ 2.2 Discriminative Vision-Language Models
The prevalent approach for training Discriminative VLMs follows the two-tower contrastive approach pioneered by CLIP [36], whereby an image and text encoder are trained on web-collected image-text pairs to learn a joint multi-modal (  i.e., vision and language) space. Subsequent works build upon CLIP by scaling the data [48, 37], improving the architecture using late/early interactions [29] or improving the training loss [48, 7]. Despite their remarkable zero-shot and representation learning abilities [36] such models were shown to have significant shortcomings related to limited language understanding capabilities, including: lack of compositionality understanding [25], manifesting bag of words behavior [47], struggling with spatial relations [25], being susceptible to typographical attacks [17], etc. Recent works aim to address these shortcomings by constructing synthetic hard negatives [47] or performing cross-modality attention [29]. However, the former does not inherently change the model’s behaviors and has been shown to potentially learn a series of shortcuts/artifacts [19]. Meanwhile, the latter is impractical for deployment at scale, as, due to the interactions between the encoders, each new query incurs an additional inference for every image within the set.
@@ 3 Method
The following sections detail our approach which we call VladVA: Vision-Language Adaptation for Discriminative Visual Assistant.
Figure 1: Overall VladVA framework: a generative LVLM is adapted into a discriminative model. At test time, the vision features ($\mathbf{f}_{v}$) are produced by passing the image $\mathbf{x}_{v}$ alongside the handcrafted and/or soft prompt $\mathbf{x}^{v}_{p}$ to the LVLM. The last token of this sequence will contain the summarized representation. Analogously, the textual features ($\mathbf{f}_{t}$) are produced by passing the prompt corresponding to the language modality ($\mathbf{x}_{p}^{t}$) through the LLM of the LVLM (i.e. the weights of the LLM are fully shared) alongside the short caption ($\mathbf{x}_{q}^{short}$). During training, the LVLM will also take as input a long, detailed caption $\mathbf{x_{q}^{long}}$.
@@ 3.1 LVLMs as zero-shot discriminative models
Despite being solely trained with an autoregressive next-token prediction loss on limited amounts of data ($<5$M), such models can act as multi-modal discriminative models in a zero-shot manner [23]. To elicit this capability, the image embedding $\mathbf{f}_{v}=\Phi_{t}(g(\Phi_{v}(\mathbf{x_{v}}),\mathbf{x}_{p})[eos]$ is obtained by passing the image alongside a handcrafted prompt $\mathbf{x}_{p}$ through the LVLM and taking the last token, prior to any generation, as the reference feature. Analogously, the text embedding $\mathbf{f}_{t}=\Phi_{t}(\mathbf{x}_{p},\mathbf{x}_{q})[eos]$ is produced by passing the handcrafted prompt $\mathbf{x}_{p}$ and input $\mathbf{x}_{q}$ query through the LLM of the LVLM and taking again the features of the last token, prior to any autoregressive step. We will refer to these particular tokens as “summary tokens”. Note that, typically, the respective handcrafted prompts for the image and text modalities are different. Finally, the similarity between an image and a given text can be computed by taking the cosine similarity between the two: $s=\texttt{cos\_sim}(\mathbf{f}_{v},\mathbf{f}_{t})$.
@@
Figure 3: Cumulative variance of the image and text embedding matrices over a set of 50 prompts on Flickr30k. Embeddings that capture more information about the input translate into a cumulative variance that requires more principal components to be explained, i.e. a higher rank embedding matrix.
Figure 4: Top-k next-to-be-predicted tokens before and after the proposed fine-tuning approach. On the right, we show the output probability distribution for each case. We observe that, when using the optimal prompt, the representations of the next token can encode diverse and more discriminative information, making potentially better-quality embeddings. This behavior is further improved by the proposed fine-tuning method.
@@
To answer these questions, we construct a testbed consisting of 1,000 image-caption pairs from Flickr30k [46], which we then use to evaluate the quality of various prompts. The prompts (50 image-text pairs in total) are constructed using ChatGPT. Each prompt pair is fed, alongside an image and its respective caption, through the LLaVA-1.5-7B model. For each image-prompt pair and, respectively, caption-prompt pair, we take its token embedding at the output position and its output probability distribution over the vocabulary and use them to compute two measures for each prompt: the average entropy of its output distributions and the cumulative variance of its embeddings. As Figs. 2 and 4 show, when the model is prompted with sentences consisting of certain keywords, such as in a few words or in one word, the model is pushed to condense the information of the image or text in the next token, resulting in an output distribution with high entropy. More importantly, when investigating the generated embeddings, we observe that higher entropy prompts result in embeddings with more spread-out cumulative variance, i.e. requiring more principal components to capture the same amount of variance, indicating an embedding matrix with a high rank (see Fig. 3). This translates into discriminative embeddings that can capture more information about the inputs, making them suitable for embedding tasks. The benefit of this behavior is illustrated in Fig. 5, which shows a positive correlation between prompts with high entropy scores and the model’s zero-shot retrieval performance. Hence, our approach should seek to produce embeddings with a) spread-out variance and b) probability distributions over the vocabulary with increased entropy.
@@ 3.2 Discriminative fine-tuning of LVLMs: from generation to discrimination
Despite exhibiting surprising innate zero-shot abilities, LVLM’s direct discriminative performance lags behind that of state-of-the-art contrastively trained VLMs. Hence, carefully designed frameworks are needed to unlock the full potential of such models. This is the very goal of our work: to introduce a well-grounded image-text adaptation/training framework that surfaces the discriminative vision-language capabilities of a generative LVLM.
Notably, our findings contradict those of the very recent (concurrent) work of [23], which found that contrastive image-text fine-tuning is detrimental and limits training to text-text contrastive learning alone. This highlights the importance of our proposed approach, which overcomes such impediments and significantly boosts the discriminative performance of the model.
@@
Data strategy: We argue for the importance of data diversity in terms of granularity and group captions according to their length: short captions ($<30$ tokens) and long captions ($30-500$ tokens). The short captions capture coarse details and summarize image content, which teaches the model to discriminate with regard to the high-level information of an image. Longer captions capture finer details in images and promote a better understanding of language concepts such as spatial relationships and compositionality. For a strong discriminative model, both are necessary. Therefore, for images missing either caption type, we use a BLIP2 [29] captioner to generate short captions and ShareGPT-4V [8] to generate long captions. This allows us to leverage both supervisory signals for training.
Training strategy: As we demonstrate in this work, the variable length of the training data poses its own challenges: unlike the case of short captions, where direct training using the well-studied contrastive loss performs well, it collapses for longer captions. This brings us to the proposed training strategy, whereby, to address this challenge, we propose a well-motivated hybrid training approach that combines a contrastive loss (see Sec. 3.2.1) and a next-token prediction loss for discriminative adaptation (see Sec. 3.2.2). Finally, as full model fine-tuning is computationally expensive, in Sec. 3.3, we detail a fine-tuning strategy that combines adapters with soft prompting.
@@ 3.2.1 Image-text contrastive alignment
During training, the contrastive loss is applied to the very same tokens used for the zero-shot evaluation, as they represent the optimal starting point for further fine-tuning. We note that the contrastive loss is mostly suitable for training using short captions ($<30$ tokens), like the ones typically used for CLIP pre-training. We found that training the model using a contrastive loss on longer captions proves challenging. Hence, to address this, in the following section, we study and propose a new formulation that enables discriminative training on variable-length data.
@@ 3.2.2 Autoregressive training for learning discriminative LVLM representations
$\mathcal{L}_{CE}=\sum_{i=1}^{L}\log p_{\theta}(u_{i}|\textbf{x}_{v},\textbf{x}_{p},\textbf{x}_{q,<i}),\vspace{-0.2cm}$ (2)
where $L$ is the length of the long caption, $\mathbf{x}_{q}$, $\mathbf{x}_{v}$ the input image and $\mathbf{x}_{p}$ the prompt, and $p_{\theta}$ as the next-token probability distribution produced by the model at input token $u_{i}$.
Intuitively, this formulation possesses multiple advantages: (1) It allows the model to learn from long captions, as predicting each and every token correctly is a challenging task (as opposed to applying the contrastive loss to long captions). (2) The decoding process encourages the condensation of information into the starting token used as a feature embedding. (3) It offers an avenue for retaining the generative capabilities of the model while strengthening its discriminative abilities.
@@ 3.2.3 Overall training loss
We apply the next-token prediction loss over the long captions and the contrastive loss over the short ones in a unified manner. During training, the templates presented to the LVLM for the image and, respectively, text modality take the following form:
@@
ASSISTANT: <long_caption> <out_token>
@@
with the contrastive loss applied between the <out_token> belonging to the image modality on one side, and the text modality on the other. Concomitantly, the next-token prediction cross-entropy loss is applied over the tokens belonging to the <long_caption>. Note that the short caption must be sufficiently different from the long caption to prevent shortcuts during training, a property that does not naturally emerge in our case due to the difference in length and annotation procedure.
@@
3.3 Parameter Efficient Adaptation
As direct fine-tuning of the LVLM is costly, especially while attempting to maintain a reasonably large batch size for contrastive learning, herein we explore and adopt soft-prompting combined with LoRA adapters, both trained under the same loss formulation of Sec. 3.2.
Soft prompting was recently proposed as an efficient task-adaptation approach for both LLM [30] and CLIP [49, 6] models, representing a direct departure from the prompt hand-crafting solution. Specifically, for a given input modality, i.e. image and text, we define a set of $n$ modality($m$)-specific learnable vectors $[\mathbf{v}_{1}^{m},\mathbf{v}_{2}^{m},\cdots,\mathbf{v}_{n}^{m}]$, $\mathbf{v}_{i}^{m}\in\mathbb{R}^{C}$ with $C$ denoting the model’s vocabulary embedding size. These vectors can be inserted across the input sequence to adjust the model’s behavior. In practice, we opt to replace the tokens belonging to the hard prompts (i.e. $\mathbf{x}_{p}$; see Sec. 3.1) with the learnable vectors, initializing their values with the embeddings of the handcrafted ones.

### v3

# VladVA: Discriminative Fine-tuning of LVLMs

@@ Abstract
Contrastively-trained Vision-Language Models (VLMs) like CLIP have become the de facto approach for discriminative vision-language representation learning. However, these models have limited language understanding, often exhibiting a “bag of words” behavior. At the same time, Large Vision-Language Models (LVLMs), which combine vision encoders with LLMs, have been shown to be capable of detailed vision-language reasoning, yet their autoregressive nature renders them less suitable for discriminative tasks.
@@ 1 Introduction
Concomitantly, inspired by the success of LLMs [5, 50] in acting as generalist assistants [16], a series of works combine pretrained vision encoders and LLMs [29, 30, 58] to construct Large Vision-Language Models (LVLMs) capable of performing interactive multi-modal conversations. Among others, these models have been shown capable of exhibiting strong reasoning and vision-language understanding capabilities, offering fine-grained and detailed responses [30, 29, 10, 12]. However, they are trained with a next-token prediction loss in an autoregressive manner, which appears less suitable for direct utilization in discriminative image-text tasks (e.g. image-text retrieval).
@@
In our approach, following the (independent) two-towers paradigm, the vision embeddings are produced by passing the image through the entire LVLM, and the text embeddings by passing the text through the LLM of the LVLM. Intuitively, for the vision embedding, the LLM acts as an information processor that refines the visual information while simultaneously aligning it with the textual representations. We coin our approach VladVA: Vision-Language Adaptation for Discriminative Visual Assistant. Our main contributions are:
  • We devise a carefully designed optimization framework that utilizes image-text pairs of variable length and granularity for model training (i.e. both short and long captions). Using this data, the model is trained with both contrastive and next-token prediction losses, which are both shown to be necessary for unlocking strong discrimination and compositionality capabilities. Our design choices are accompanied by ablation studies, which justify the necessity of our framework’s components.
@@ 2.1 Large Vision Language Models (LVLMs)
Despite their strong generative and comprehension abilities [37], current LVLMs are primarily restricted to generative tasks. Only very recently, Jiang et al. [24], inspired by the recent progress in NLP [4, 27] adapted a LLaVA-NeXT [29] model to discriminative tasks using a contrastive-like loss and text data only. We note that unlike [24], we introduce a training framework that learns from multi-turn image-text pairs (as opposed to text only) using a novel formulation that jointly combines a contrastive loss with a next-token prediction, reflecting the data characteristics and inducing a gradual representation buildup. Concurrently, VLM2Vec [25] adapts an LVLM for multi-modal retrieval. However, it uses a different loss and training strategy (no generative loss, no short-long captions training, no soft prompting). We compare our approach with both E5-V and VLM2Vec, significantly improving upon their results despite using smaller/lighter models.
@@ 2.2 Discriminative Vision-Language Models
The prevalent approach for training Discriminative VLMs follows the two-tower contrastive approach pioneered by CLIP [42], whereby an image and text encoder are trained on web-collected image-text pairs to learn a joint multi-modal (i.e. vision and language) space. Subsequent works build upon CLIP by scaling the data [56, 44, 7], improving the architecture using late/early interactions [32] or improving the training loss [56, 8]. Despite their remarkable zero-shot and representation learning abilities [42] such models were shown to have significant shortcomings related to limited language understanding capabilities, including: lack of compositionality understanding [28], manifesting bag of words behavior [55], struggling with spatial relations [28], being susceptible to typographical attacks [18], etc. Recent works aim to address these shortcomings by constructing synthetic hard negatives [55] or performing cross-modality attention [32]. However, the former does not inherently change the model’s behaviors and has been shown to potentially learn a series of shortcuts/artifacts [20]. Meanwhile, the latter is impractical for deployment at scale, as, due to the interactions between the encoders, each new query incurs an additional inference for every image within the set.
@@ 3 Method
Herein, we present VladVA (Vision-Language Adaptation for Discriminative Visual Assistant), our novel approach for discriminative fine-tuning of LVLMs that results in strong discriminative and compositional capabilities. This section is structured as follows: Sec. 3.1 briefly introduces the architecture, detailing how LVLMs can be used as discriminators in a zero-shot manner. Sec. 3.2 details the core component of our approach: a carefully designed optimization framework that utilizes image-text pairs of variable length and granularity for training the model with both contrastive (Sec. 3.2.1) and next-token prediction (Sec. 3.2.2) losses, showcasing that contrastive is best with short captions and autoregressive with long captions. In Sec. 3.3, we present our parameter-efficient adaptation, while Sec. 3.4 analyzes how the model’s behavior changes after training.
Figure 1: Overall VladVA framework: a generative LVLM is adapted into a discriminative model with the help of (1) a contrastive training loss (Sec. 3.2.1), and (2) an autoregressive loss (Sec. 3.2.2). The first one is applied on image-text pairs with short(er) captions, encouraging the last token produced by both modalities to be discriminative. The second one, jointly optimized with the first one, is applied only on longer captions and allows the model to learn fine-grained details.
@@ 3.1 LVLMs as zero-shot discriminative models
Despite being solely trained with an autoregressive next-token prediction loss on limited amounts of data ($<5$M), such models can act as multi-modal discriminative models in a zero-shot manner [24]. To elicit this capability, the image embedding $\mathbf{f}_{v}=\Phi_{t}(g(\Phi_{v}(\mathbf{x}_{v}),\mathbf{x}^{v}_{p}))[eos]$ is obtained by passing the image alongside a handcrafted image prompt $\mathbf{x}^{v}_{p}$ (e.g., “in one word, describe the image”) through the LVLM and taking the output representation of the last token. Analogously, the text embedding $\mathbf{f}_{t}=\Phi_{t}(\mathbf{x}^{t}_{p},\mathbf{x}_{q})[eos]$ is produced by passing the handcrafted text prompt $\mathbf{x}^{t}_{p}$ (e.g., “in one word, describe the text”) and input query $\mathbf{x}_{q}$ through the LLM (of the LVLM) and taking again the output representation of the last token. We will refer to these particular tokens as “summary tokens” (summarizing image and text information, respectively). Note that, typically, the respective handcrafted prompts for the image ($\mathbf{x}^{v}_{p}$) and text ($\mathbf{x}^{t}_{p}$) modalities are different. Finally, the similarity between an image and a text query can be computed by taking the cosine similarity between the two: $s=\texttt{cos\_sim}(\mathbf{f}_{v},\mathbf{f}_{t})$.
@@
Figure 3: Cumulative variance of the image and text embedding matrices over a set of 50 prompts on Flickr30k. Embeddings that capture more information about the input translate into a cumulative variance that requires more principal components to be explained, i.e. a higher-rank embedding matrix.
Figure 4: Top-k next-to-be-predicted tokens before and after VladVA fine-tuning (our approach). On the right, we show the output probability distribution for each case. When using the best prompt (“Summarize the provided image in one word”), the representations of the next token can encode diverse and more discriminative information, making potentially better-quality embeddings. This behavior is further improved after VladVA fine-tuning.
@@
To answer these questions, we construct a testbed consisting of 1,000 image-caption pairs from Flickr30k [54], which we then use to evaluate the quality of various prompts. The prompts (50 image-text pairs in total) are constructed using ChatGPT. Each prompt pair is fed, alongside an image and its respective caption, through the LLaVA-1.5-7B model. For each image-prompt pair and caption-prompt pair, we extract the token embedding at the output position and the corresponding output probability distribution over the vocabulary. These are then used to compute two metrics for each prompt: the average entropy of its output distributions and the cumulative variance of its embeddings. As Figs. 2 and 4 show, when the model is prompted with sentences consisting of specific keywords, such as in a few words or in one word, the model is pushed to condense the information of the image or text in the next token, resulting in an output distribution with high entropy. More importantly, when investigating the generated embeddings, we observe that higher entropy prompts result in embeddings with more spread-out cumulative variance, i.e. requiring more principal components to capture the same amount of variance, indicating an embedding matrix with a high rank (see Fig. 3). This translates into discriminative embeddings that can capture more information about the inputs, making them suitable for embedding tasks. The benefit of this behavior is illustrated in Fig. 5, which shows a positive correlation between prompts with high entropy scores and the model’s zero-shot retrieval performance. Hence, our approach should seek to produce embeddings with a) spread-out variance and b) probability distributions over the vocabulary with increased entropy.
@@ 3.2 Discriminative fine-tuning of LVLMs: from generation to discrimination
Despite exhibiting surprising innate zero-shot abilities, LVLM’s direct discriminative performance lags behind that of state-of-the-art contrastively trained VLMs. Hence, carefully designed frameworks are needed to unlock the full potential of such models. This is the very goal of our work: to introduce a well-grounded adaptation/training framework that surfaces the discriminative image-text capabilities of a generative LVLM.
Notably, our findings contradict those of the very recent work of [24], which found that contrastive image-text fine-tuning is detrimental and limits training to text-text contrastive learning alone. This highlights the importance of our proposed approach, which overcomes such impediments and significantly boosts the discriminative performance of the model.
@@
Data strategy: We argue for the importance of data diversity in terms of granularity and group captions according to their length: short captions ($<30$ tokens) and long captions ($30-500$ tokens). The short captions capture coarse details and summarize image content teaching the model to discriminate with regard to high-level image information. Longer captions capture finer image details and promote a better understanding of language concepts such as spatial relationships and compositionality. For a strong discriminative model, both are necessary. Therefore, for images missing either caption type, we use a BLIP2 [32] captioner to generate short captions and ShareGPT-4V [9] to generate long captions. This allows us to leverage both supervisory signals for training.
Training strategy: As we demonstrate in this work, the variable length of the training data poses its own challenges: unlike the case of short captions, where training using the well-studied contrastive loss performs well, it collapses for longer captions. This brings us to the proposed training strategy, whereby, to address this challenge, we propose a hybrid training approach that combines a contrastive loss (see Sec. 3.2.1) and a next-token prediction loss for discriminative adaptation (see Sec. 3.2.2). Finally, as full model fine-tuning is computationally expensive, in Sec. 3.3, we detail a fine-tuning strategy that combines adapters with soft prompting.
@@ 3.2.1 Image-text contrastive alignment
During training, the contrastive loss is applied to the very same tokens used for the zero-shot evaluation, as they represent the optimal starting point for further fine-tuning (Sec. 3.1). We note that the contrastive loss is mostly suitable for training using short captions $\mathbf{x}^{short}_{q}$ (i.e. $<30$ tokens), like the ones typically used for CLIP pre-training. We found that training the model using a contrastive loss on longer captions proves challenging. Hence, to address this, in the following section, we study and propose a new formulation that enables discriminative training on variable-length data.
@@ 3.2.2 Autoregressive training for learning discriminative LVLM representations
$\mathcal{L}_{CE}=\sum_{i=1}^{L}\log p_{\theta}(u_{i}|\textbf{x}_{v},\textbf{x}^{v}_{p},\textbf{x}^{long}_{q,<i}),\vspace{-0.2cm}$ (2)
where $L$ is the length of the long caption $\mathbf{x}^{long}_{q}$, $\mathbf{x}_{v}$ the input image, and $\mathbf{x}_{p}^{v}$ the prompt which prompts the model to describe the image in detail (e.g., “Describe the image in detail”), and $p_{\theta}$ the next-token probability distribution learned by the model.
Intuitively, this formulation possesses multiple advantages: (1) It allows the model to learn from long captions, as predicting each and every token correctly is a challenging task (as opposed to applying the contrastive loss to long captions); (2) The decoding process encourages the condensation of information into the starting token used as a feature embedding; (3) It offers an avenue for retaining the generative capabilities of the model while strengthening its discriminative abilities.
@@ 3.2.3 Overall training loss
As depicted in Fig. 1, we apply the next-token prediction loss over the long captions and the contrastive loss over the short ones in a unified manner. During training, the templates presented to the LVLM for the image and text modality take the following form:
@@
ASSISTANT: <long_caption>
@@
with the contrastive loss applied on the output representations <out_token> for the image modality and <out_token> for the text modality. Concomitantly, the next-token prediction loss is applied on the tokens of the <long_caption>. Generally, the short caption must be sufficiently different from the long caption to prevent shortcuts during training, a property that naturally emerges in our case due to the difference in length and annotation procedure. Note that the distinction between long and short captions is made only during training. At test time, the model is used in discriminative mode as detailed in Sec. 3.1.
@@
3.3 Parameter-efficient adaptation
As direct fine-tuning of the LVLM is costly, especially when maintaining a reasonably large batch size for contrastive learning, herein, we adopt parameter-efficient training with soft-prompting combined with LoRA adapters, both trained under the same loss formulation of Sec. 3.2.
Soft prompting was recently proposed as an efficient task-adaptation approach for both LLM [33] and CLIP [57, 6] models, representing a direct departure from the prompt hand-crafting solution. Specifically, for a given input modality, i.e. image and text, we define a set of $n$ modality($m$)-specific learnable vectors $[\mathbf{v}_{1}^{m},\mathbf{v}_{2}^{m},\cdots,\mathbf{v}_{n}^{m}]$, $\mathbf{v}_{i}^{m}\in\mathbb{R}^{C}$ with $C$ denoting the model’s vocabulary embedding size. These vectors can be inserted across the input sequence to adjust the model’s behavior. In practice, we opt to replace the tokens belonging to the hard prompts (i.e. $\mathbf{x}_{p}^{v}$ and $\mathbf{x}_{p}^{t}$; see Sec. 3.1) with the learnable vectors, initializing their values with the embeddings of the handcrafted ones.